# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Appendix B Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
This appendix explains why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as practical baselines in the high-dimensional uncertain markets targeted in this paper. DP is conceptually sound in low-dimensional Markovian settings (Fleming and Soner, 2006; Pham, 2009), but two obstructions dominate in our regime: (i) value-function approaches require accurate derivatives (often including mixed sensitivities) to recover optimal portfolios, and this becomes prohibitive as dimension and nonlinearity grow; (ii) principled parameter uncertainty magnifies both the state dimension and the aggregation difficulties. These points are standard for grid-based HJB solvers (Bellman, 1961; Kushner and Dupuis, 2001) and remain relevant for modern deep PDE surrogates (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), especially under partial information/belief-state formulations (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
@@
B.1 The value-function target for latent $\theta$ is a belief-state functional
@@
$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (86)
where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{blind}}$ denotes admissible controls measurable with respect to the observable filtration (so the deployed policy cannot take $\theta$ as an input).
Why $(t,X_{t},Y_{t})$ is not a closed DP state.
In general, the observable history carries information about $\theta$, so the continuation value at time $t$ depends on the conditional law
@@
not only on $(t,X_{t},Y_{t})$. This is the standard partially observed control phenomenon: DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
@@
The principled DP value function is therefore a functional of a probability measure:
$V(t,x,y,\mu):=\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}^{t,x,y,\mu}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\mu\in\mathcal{P}(\Theta),$ (88)
so the associated HJB equation is posed on a space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Under our inference-agnostic interface, $q$ need not belong to a conjugate family and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional class. Consequently, a DP/PDE baseline would be forced to approximate $V$ over an infinite-dimensional belief state, which is not the computational target of this paper.
B.2 Even with deterministic parameters, HJB solvers face a derivative bottleneck in high dimension
Classical HJB: curse of dimensionality and full nonlinearity.
With deterministic parameters, DP leads to an HJB for $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009). Grid-based solvers scale exponentially in the state dimension (Bellman, 1961; Kushner and Dupuis, 2001). In portfolio problems with $d$ assets and $m$ factors, the natural state already has dimension $m+2$, so even modest discretizations require $N^{m+2}$ grid points.
Moreover, realistic features (constraints, transaction costs, non-affine dynamics) typically yield fully nonlinear HJBs. Recovering the optimal portfolio from a value function also requires derivatives (schematically, in Merton-type problems)
@@ B.2 Even with deterministic parameters, HJB solvers face a derivative bottleneck in high dimension
and in factor-driven settings intertemporal hedging demands are driven by mixed sensitivities (e.g. $V_{xy}$). Thus the numerical burden is not only approximating $V$ but also obtaining stable first/second/mixed derivatives in high dimension.
B.3 Deep PDE surrogates remove grids but not the value-function derivative bottleneck
PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019), but for fully nonlinear portfolio HJBs they remain value-function-based: they must implicitly learn high-dimensional gradients/Hessians and, crucially, mixed sensitivities that drive intertemporal hedging. In practice this can induce nonconvex and ill-conditioned training objectives (due to control suprema and nonlinear derivative dependence) and training signals that do not reliably control the specific derivative components needed for stable high-dimensional portfolio policies.
This is not a claim that deep PDE methods are “invalid”; rather, under our high-dimensional and $\theta$-blind interface, they are not practical baselines relative to direct, simulation-only policy optimization methods.
B.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
A natural alternative is to condition on $\theta$ and solve a family of full-information problems.
$\theta$-conditional HJB.
@@ B.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
for the controlled generator $\mathcal{L}^{\pi,\theta}$ (Fleming and Soner, 2006; Pham, 2009). This yields a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$, which is infeasible in our deployable $\theta$-blind setting.
Supremum–expectation non-commutativity.
More fundamentally, even perfect knowledge of the $\theta$-conditional optimizer does not solve the $\theta$-blind objective, because
$\sup_{\pi\in\mathcal{A}^{\mathrm{blind}}}\mathbb{E}_{\theta\sim q}\!\left[J(\pi,\theta)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J(\pi,\theta)\right],\hskip 18.49988ptJ(\pi,\theta):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (91)
and the inequality is typically strict. The right-hand side corresponds to the (infeasible) $\theta$-adaptive controller.
B.5 Why “sample $\theta$ only in the PDE loss” is mismatched to the ex–ante objective
One might try to train a single value network $V_{\varphi}(t,x,y)$ that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a collocation/pathwise residual loss:
@@ B.5 Why “sample $\theta$ only in the PDE loss” is mismatched to the ex–ante objective
where $\mathrm{HJB}_{\theta}(\cdot)$ is the full-information HJB operator for parameter $\theta$. This is appealing but generally mismatched to (86):
  • Different $\theta$ impose different (often conflicting) optimality conditions at the same $(t,x,y)$, so there is generally no single $V$ satisfying $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$ simultaneously.
  • Minimizing an expected squared PDE residual (92) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls.
  • Portfolio feedback rules are typically extracted through nonlinear operations on derivatives (ratios, inverses), so small value-function approximation errors can induce large control errors.
B.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
The aggregation issue is visible even in a one-asset Merton model with static latent drift. Let
@@ B.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
and restrict to constant fractions $\pi$ over a remaining horizon $\tau$. For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, the $\theta$-conditional full-information constant Merton rule is
@@
Averaging this infeasible rule gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
@@
which exhibits shrinkage due to estimation uncertainty ($p>0$) and depends on the horizon $\tau$. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $p>0$.
B.7 Implication for baselines in this paper
The discussion above motivates our baseline choices in the main text:
  • We validate on controlled drift/premium benchmarks where closed-form decision-time Gaussian targets are available and match the fixed-$q$ deployability restriction.
  • For learning-based comparisons, we emphasize simulation-only direct policy search baselines (e.g., PPO) that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
Appendix C Proof of Theorem 2
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see the main BPTT–BSDE correspondence result and proof in Huh et al. (2025b)). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.  (This appendix replaces the current Appendix C in the draft.)
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (45)–(46). It does not use the $\theta$-blind $q$-aggregated stationarity condition (Theorem 1 in Section 3.2). Those constructions affect only stage 2 projection targets and are irrelevant to the BPTT–PMP convergence itself.
@@ Appendix C Proof of Theorem 2
$\mathbb{G}^{\theta}:=(G_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988ptG_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and $L^{2}$ projections below are taken with respect to $G_{t_{k}}^{\theta}$.
@@
Key point. The raw BPTT “adjoints” are pathwise derivatives of the terminal utility along the full discrete computation graph. Hence, for $k<N$, they are typically $G_{T}^{\theta}$-measurable and need not be $G_{t_{k}}^{\theta}$-measurable (they can depend on future increments). In contrast, the stochastic maximum principle (SMP/PMP) costate is $\mathbb{G}^{\theta}$-adapted. Therefore, the correspondence is formulated after a standard one-step conditional $L^{2}$ projection (equivalently, a discrete-time BSDE representation), which turns pathwise objects into adapted discrete adjoints.
@@
Let $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\theta},Y_{k}^{\theta})$, where $\pi_{\varphi}$ is $\theta$-blind but evaluated along the $\theta$-conditional trajectory.
@@
Step 1: Conditioning on $\theta$ and uniformity of bounds. Fix a compact set $K\subset\Theta$. Assume the coefficients in (45)–(46) satisfy the usual Lipschitz and linear-growth conditions uniformly over $\theta\in K$, and that the block covariance structure of $(W,W^{Y})$ (including instantaneous correlation) is uniformly nondegenerate on $K$. Assume also the regularity/integrability needed for the $\theta$-conditional SMP (including the blocks relevant for the portfolio Hamiltonian gradient) to be well posed. Then, for each fixed $\theta\in K$, the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{K}\,\Delta t^{1/2},$
with a constant $C_{K}$ that can be chosen independently of $\theta\in K$. These are exactly the deterministic assumptions used in Huh et al. (2025b), now stated uniformly on $K$.
@@
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Under fixed $\theta$, consider the Euler scheme on the grid $(t_{k})$:
$Y_{k+1}^{\theta}=Y_{k}^{\theta}+a(Y_{k}^{\theta},\theta)\Delta t+\beta(Y_{k}^{\theta},\theta)\Delta W_{k}^{Y},$
$X_{k+1}^{\theta}=X_{k}^{\theta}+X_{k}^{\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\theta},\theta)\Big)\Delta t+X_{k}^{\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\theta},\theta)\Delta W_{k},$
with terminal reward $U(X_{N}^{\theta})$. Define the discrete pathwise wealth costate
$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\theta}}\,U(X_{N}^{\theta}),\hskip 18.49988ptk=0,\dots,N,$
@@
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\theta}}.$
@@
Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph. The one-step algebraic form of the backward recursion coincides with the deterministic-parameter analysis in Huh et al. (2025b), with the replacements
$\mu\mapsto b(\cdot,\theta),\hskip 18.49988pt\sigma\mapsto\sigma(\cdot,\theta),$
and with the factor block handled exactly as in the wealth–factor extension therein. All one-step remainder terms are controlled by standard Taylor/Euler estimates with constants uniform in $\theta\in K$.
@@
Step 3: Conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables (fixed $\theta$). Fix $\theta\in K$. Let $\Delta W_{k}:=W_{t_{k+1}}-W_{t_{k}}$ and $\Delta W_{k}^{Y}:=W^{Y}_{t_{k+1}}-W^{Y}_{t_{k}}$. Uniform nondegeneracy of the block covariance of $(\Delta W_{k},\Delta W_{k}^{Y})$ implies that the conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined (and yields unique coefficients).
(a) One-step projection (defining the discrete martingale coefficients). Define $(z_{k}^{\theta},\tilde{z}_{k}^{\theta})$ by the conditional orthogonality relations (equivalently, the $L^{2}$-projection) such that
@@
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid G_{t_{k}}^{\theta}\right]+z_{k}^{\theta}\Delta W_{k}+\tilde{z}_{k}^{\theta}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
@@
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $G_{t_{k}}^{\theta}$.
@@
(b) Adapted discrete adjoints (the objects compared to PMP). Define the adapted discrete-time costates by conditional expectation:

### v3

@@ Appendix B Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Appendix C Why dynamic programming and deep PDE surrogates are not primary baselines under the $\theta$-blind fixed-$q$ interface
This appendix clarifies why we do not treat classical dynamic programming (DP/HJB) or value-function-based deep PDE surrogates (PINNs / deep BSDE methods) as primary baselines for the high-dimensional uncertain markets targeted in this paper. The main obstruction is not that PDE solvers are “invalid”; rather, under our $\theta$-blind fixed-$q$ deployability interface, the natural DP/PDE/BSDE target is typically not a low-dimensional Markov value function. Derivative instability and high-dimensional conditioning issues are secondary to this target mismatch.
@@
C.1 The value-function target for latent $\theta$ is a belief-state functional
@@
$\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}J(\pi)\;:=\;\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}_{\theta\sim q}\!\left[\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right]\right],$ (91)
where $\theta\sim q$ is latent, fixed along each episode, and $\mathcal{A}^{\mathrm{obs}}$ denotes the class of deployable controls, i.e. controls measurable w.r.t. the observable filtration (so deployed policies cannot condition on $\theta$). (One may further restrict to the feedback subclass $\mathcal{A}^{\mathrm{fb}}$ as defined in Section 2.1.)
Why $(t,X_{t},Y_{t})$ is not a closed DP state under latent $\theta$.
Even if the prior $q$ is fixed at decision time (and even if the controller is constrained to be $\theta$-blind), the conditional law of $\theta$ given observations typically evolves as data arrive. Formally, the continuation value at time $t$ depends on the conditional law
@@
not only on $(t,X_{t},Y_{t})$. This is the standard partially observed control phenomenon: the unconditional (mixed-over-$\theta$) law of the observed state is generally not Markov in $(X_{t},Y_{t})$, and DP becomes Markov only after augmenting the state by a belief (or a finite-dimensional sufficient statistic when available) (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
@@
The principled value object is therefore a functional of a probability measure:
$V(t,x,y,\eta):=\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}^{t,x,y,\eta}\!\left[U(X_{T}^{\pi})\right],\hskip 18.49988pt\eta\in\mathcal{P}(\Theta),$ (93)
where $\mathbb{E}^{t,x,y,\eta}$ denotes the expectation for a system started from $(t,x,y)$ with latent parameter $\theta\sim\eta$. Hence the associated DP/HJB is posed on a (typically infinite-dimensional) space of measures in general (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
Under our inference-agnostic interface, $q$ need not be conjugate and we do not specify a filtering model that would keep $q_{t}$ in a tractable finite-dimensional family. Consequently, any value-function baseline that assumes a low-dimensional Markov target is, by construction, solving a different object from the $\theta$-blind fixed-$q$ problem studied in this paper.
C.2 Even with deterministic parameters, value-function baselines are indirect under our interface
When parameters are deterministic, DP leads to an HJB for a Markov value function $V(t,x,y)$ (Fleming and Soner, 2006; Pham, 2009). In portfolio problems, however, optimal feedback rules are typically recovered through nonlinear operations on derivatives of $V$, e.g.
@@ C.2 Even with deterministic parameters, value-function baselines are indirect under our interface
with intertemporal hedging driven by mixed sensitivities such as $V_{xy}$. This “derivative bottleneck” is a practical issue for grid-based solvers (Bellman, 1961; Kushner and Dupuis, 2001) and also affects neural surrogates, but it is not the core reason we exclude DP/PDE baselines here. The core reason is the belief-state mismatch in Section C.1.
C.3 Deep PDE surrogates (PINNs / deep BSDE) do not resolve the target mismatch
PINNs and deep BSDE methods replace grids with neural approximators trained on sampled points/paths (Raissi et al., 2019; Sirignano and Spiliopoulos, 2018; Han et al., 2018; Beck et al., 2019; Huré et al., 2020). However, both remain value-function-based PDE/BSDE solvers: they implicitly assume a (finite-dimensional) Markov target for the value/decoupling field. Under latent $\theta$, a solver must therefore choose between:
  • learning $\theta$-conditional objects (making the policy/value depend on $\theta$, hence not deployable); or
  • enforcing a $\theta$-agnostic Markov representation (e.g. a single $V(t,x,y)$) by sampling $\theta$ only inside the fitting loss, which generally does not correspond to the $\theta$-blind ex–ante optimum (91).
The next two subsections make this mismatch explicit.
C.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
$\theta$-conditional HJB / BSDE.
@@ C.4 Why $\theta$-conditional value functions do not solve the $\theta$-blind ex–ante problem
(Fleming and Soner, 2006; Pham, 2009), yielding a $\theta$-dependent optimizer $\pi^{\star}(t,x,y,\theta)$. This is infeasible under our $\theta$-blind deployability restriction.
Supremum–expectation non-commutativity is the structural obstruction.
Even perfect access to $\theta$-conditional optimizers does not solve the $\theta$-blind objective because
$\sup_{\pi\in\mathcal{A}^{\mathrm{obs}}}\mathbb{E}_{\theta\sim q}\!\left[J^{\theta}(\pi)\right]\;\leq\;\mathbb{E}_{\theta\sim q}\!\left[\sup_{\pi}J^{\theta}(\pi)\right],\hskip 18.49988ptJ^{\theta}(\pi):=\mathbb{E}\!\left[U(X_{T}^{\pi})\mid\theta\right],$ (96)
and the inequality is typically strict. The right-hand side corresponds to an (infeasible) $\theta$-adaptive controller. Equivalently, our Pontryagin stationarity is $q$-aggregated over $\theta$-blind variations, not enforced parameter-by-parameter.
C.5 Why “sampling $\theta$ only in the PDE/BSDE loss” is mismatched to the ex–ante objective
A common workaround is to learn a single value/decoupling network that does not take $\theta$ as input, and sample $\theta\sim q$ only inside a training loss. For PINNs this appears as an averaged collocation residual; for deep BSDE methods it appears as an averaged pathwise BSDE/HJB fitting objective. Schematically,
@@ C.5 Why “sampling $\theta$ only in the PDE/BSDE loss” is mismatched to the ex–ante objective
where $\mathrm{HJB}_{\theta}(\cdot)$ denotes the full-information HJB operator for parameter $\theta$.
The issue is conceptual (target mismatch), not merely numerical:
  • Conflicting optimality conditions. Different $\theta$ impose different local optimality conditions at the same $(t,x,y)$, so there is generally no single low-dimensional Markov $V$ (or decoupling field) that simultaneously satisfies $\mathrm{HJB}_{\theta}(V)=0$ for many $\theta$.
  • Loss mismatch. Minimizing an averaged PDE/BSDE fitting loss such as (97) is not equivalent to maximizing expected terminal utility under $\theta$-blind controls in (91).
  • Wrong stationarity notion. Our deployable optimum is characterized by a $q$-aggregated first-order condition with $\theta$-blind admissible variations; averaging full-information HJB/BSDE conditions across $\theta$ targets a different stationarity concept.
Therefore, while $\theta$-sampling can be a sensible approximation heuristic, it does not produce a principled DP/PDE/BSDE baseline aligned with the fixed-$q$, $\theta$-blind optimality notion studied in this paper.
C.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
The aggregation issue is visible even in a one-asset Merton model with static latent drift:
@@ C.6 A toy calculation: $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\star}$ under CRRA
and constant fractions $\pi$ over remaining horizon $\tau$. For CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>1$, the $\theta$-conditional full-information rule is
@@
Averaging gives $\mathbb{E}[\pi^{\star}(\theta)]=m/(\gamma\sigma^{2})$. By contrast, the deployable $\theta$-blind ex–ante optimizer within the constant-fraction class is
@@
which exhibits shrinkage and horizon dependence. Thus $\mathbb{E}[\pi^{\star}(\theta)]\neq\pi_{q}^{\mathrm{const}}(\tau)$ whenever $P>0$.
C.7 Implication for baselines in this paper
The discussion above motivates our baseline choices:
  • We validate on controlled decision-time Gaussian benchmarks that match the fixed-$q$ deployability restriction.
  • For learning-based comparisons, we emphasize simulation-only direct policy optimization methods that operate under the same observation restriction (no access to $\theta$; $\theta$ sampled only inside the simulator), rather than value-function PDE/BSDE solvers whose natural DP target is belief-augmented.
Appendix D Proof of Theorem 2
Proof of Theorem 2.
Theorem 2 extends the BPTT–PMP (equivalently, BPTT–BSDE) correspondence established for deterministic-parameter models in our prior work on PG–DPO (see Huh et al. (2025)). Here the only substantive change is that the market coefficients are indexed by a random but frozen parameter $\theta\sim q$, and we need convergence statements that hold conditionally on $\theta$ and uniformly over $\theta$ in compact subsets of $\Theta$.
Important remark (what this proof does not use). This proof concerns the $\theta$-conditional Pontryagin adjoint/costate for the fixed-$\theta$ control problem induced by (39)–(40). It does not use the $q$-aggregated $\theta$-blind stationarity condition (e.g. Theorem 1), which enters only the stage 2 projection target.
@@ Appendix D Proof of Theorem 2
$\mathbb{G}^{\theta}:=(\mathcal{G}_{t}^{\theta})_{t\in[0,T]},\hskip 18.49988pt\mathcal{G}_{t}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\}\big)\ \text{(with the usual augmentation)}.$
All conditional expectations and conditional $L^{2}$ projections below are taken w.r.t. $\mathcal{G}_{t_{k}}^{\theta}$.
@@
Key point. The raw BPTT “adjoints” are pathwise derivatives of the terminal utility along the full discrete computation graph. Hence, for $k<N$, they are typically $\mathcal{G}_{T}^{\theta}$-measurable and need not be $\mathcal{G}_{t_{k}}^{\theta}$-measurable (they can depend on future increments). In contrast, the stochastic maximum principle (SMP/PMP) costate is $\mathbb{G}^{\theta}$-adapted. Therefore, the correspondence is formulated after a standard one-step conditional $L^{2}$ projection (equivalently, a discrete-time BSDE representation), which turns pathwise objects into adapted discrete adjoints.
@@
Let $W$ be $d_{W}$-dimensional and $W^{Y}$ be $d_{Y}$-dimensional (allowing instantaneous correlation). Fix $\Delta t>0$, $t_{k}:=k\Delta t$, $k=0,\dots,N$, with $N\Delta t=T$. For readability we suppress the policy parameters $\varphi$ and write $\pi_{k}:=\pi_{\varphi}(t_{k},X_{k}^{\Delta t,\theta},Y_{k}^{\Delta t,\theta})$, where $\pi_{\varphi}$ is $\theta$-blind.
Standing assumptions (smooth Markov regime; uniform on compacts). Fix a compact set $\Theta_{0}\subset\Theta$. Assume the following hold uniformly for $\theta\in\Theta_{0}$:
  • Uniform well-posedness and moments. The coefficients in (39)–(40) satisfy the usual Lipschitz and linear-growth conditions (in the state variables), yielding unique strong solutions and uniform $L^{2}$ moment bounds.
  • Uniform smoothness. The coefficients are sufficiently smooth in state variables so that the $\theta$-conditional adjoint BSDE is well posed and the state-derivative blocks used in the portfolio Hamiltonian gradient exist and are square-integrable. (It is sufficient to assume $a,b,\beta,\sigma$ are $C^{2}$ in state variables with derivatives obeying uniform-in-$\theta$ growth/Lipschitz bounds on $\Theta_{0}$.)
  • Policy regularity. The deployed feedback $(t,x,y)\mapsto\pi_{\varphi}(t,x,y)$ is $C^{2}$ in $(x,y)$ on the working domain, with the required derivatives uniformly bounded there.
  • Smooth decoupling field (identifying the blocks used here). In the smooth Markov regime, the $\theta$-conditional adjoint admits a decoupling field $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, and we may identify $p_{x,t}^{\theta}:=\partial_{x}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$, $p_{y,t}^{\theta}:=\partial_{y}u^{\theta}(t,X_{t}^{\pi,\theta},Y_{t}^{\theta})$. These are precisely the sensitivity blocks needed for the portfolio Hamiltonian gradient.
  • One-step projection nondegeneracy. The conditional covariance matrix of $(\Delta W_{k},\Delta W_{k}^{Y})$ given $\mathcal{G}_{t_{k}}^{\theta}$ is uniformly nondegenerate on $\Theta_{0}$, so that the one-step conditional $L^{2}$-projection onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ is well defined with unique coefficients.
These are the deterministic assumptions used in Huh et al. (2025), now imposed conditionally on $\theta$ and uniformly over $\theta\in\Theta_{0}$.
@@
Step 1: Conditioning on $\theta$ and uniformity of forward Euler error. Under the standing assumptions, for each fixed $\theta\in\Theta_{0}$ the controlled SDE system is well posed and admits uniform-in-time $L^{2}$ moment bounds. Moreover, the Euler–Maruyama scheme enjoys the standard strong error bound
$\sup_{t\in[0,T]}\mathbb{E}\!\left[\left\|(X_{t}^{\pi,\theta},Y_{t}^{\theta})-(X_{t}^{\Delta t,\theta},Y_{t}^{\Delta t,\theta})\right\|^{2}\right]^{1/2}\leq C_{\Theta_{0}}\,\Delta t^{1/2},$
with a constant $C_{\Theta_{0}}$ independent of $\theta\in\Theta_{0}$.
@@
Step 2: Discrete forward scheme and BPTT pathwise adjoints (fixed $\theta$). Fix $\theta\in\Theta_{0}$. Consider the Euler scheme on the grid $(t_{k})$:
$Y_{k+1}^{\Delta t,\theta}=Y_{k}^{\Delta t,\theta}+a(Y_{k}^{\Delta t,\theta},\theta)\Delta t+\beta(Y_{k}^{\Delta t,\theta},\theta)\Delta W_{k}^{Y},$
$X_{k+1}^{\Delta t,\theta}=X_{k}^{\Delta t,\theta}+X_{k}^{\Delta t,\theta}\Big(r+\pi_{k}^{\top}b(Y_{k}^{\Delta t,\theta},\theta)\Big)\Delta t+X_{k}^{\Delta t,\theta}\,\pi_{k}^{\top}\sigma(Y_{k}^{\Delta t,\theta},\theta)\Delta W_{k},$
with terminal reward $U(X_{N}^{\Delta t,\theta})$.
Define the discrete pathwise wealth costate
$p_{k}^{\mathrm{pw},\theta}:=\frac{\partial}{\partial X_{k}^{\Delta t,\theta}}\,U(X_{N}^{\Delta t,\theta}),\hskip 18.49988ptk=0,\dots,N,$
@@
$p_{x,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial X_{k}^{\Delta t,\theta}},\hskip 18.49988ptp_{y,k}^{\mathrm{pw},\theta}:=\frac{\partial p_{k}^{\mathrm{pw},\theta}}{\partial Y_{k}^{\Delta t,\theta}}.$
@@
Automatic differentiation/BPTT computes $\{(p_{k}^{\mathrm{pw},\theta},p_{x,k}^{\mathrm{pw},\theta},p_{y,k}^{\mathrm{pw},\theta})\}_{k=0}^{N}$ via the backward chain rule along the discrete forward graph. Conditionally on $\theta$, this is exactly the deterministic-parameter BPTT construction of Huh et al. (2025).
@@
Step 3: One-step conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables. Fix $\theta\in\Theta_{0}$. On each step, perform the conditional $L^{2}$-projection of $p_{k+1}^{\mathrm{pw},\theta}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$. Equivalently, define regression coefficients $z_{k}^{\theta}\in\mathbb{R}^{d_{W}}$ and $\tilde{z}_{k}^{\theta}\in\mathbb{R}^{d_{Y}}$ such that
@@
$p_{k+1}^{\mathrm{pw},\theta}=\mathbb{E}\!\left[p_{k+1}^{\mathrm{pw},\theta}\mid\mathcal{G}_{t_{k}}^{\theta}\right]+(z_{k}^{\theta})^{\top}\Delta W_{k}+(\tilde{z}_{k}^{\theta})^{\top}\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
@@
where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$.
@@
Define the adapted discrete-time adjoint blocks by conditional expectation: